OpenStudy (anonymous):

Need help in creating a word problem, well correcting one that is from the following directions. A parent is looking for day care for her child one charges an hourly rate, this part is fine i have, the other charges a fixed fee per week for a certain number of hours then an hourly rate. This is what I need help with. Below is the problem i wrote

OpenStudy (anonymous):

selections Kiddie Castle which is a daycare facility about half way between home and work, the other is Cultural Au Pair in which the person comes to your home and takes care of the child there. Kiddie Castle charges a weekly rate of $112 plus $5.00 per hour of care for Joey. Cultural Au Pair charges a hourly rate of $7.00,

OpenStudy (anonymous):

before this i had kiddie castle with a flat rate of 224 for 35 hours a week then 6.oo per hour after and cultural au pair was the same. and when i wrote the equations fo rhtis one i was told my variables were wrong. i had kidde with y=224 + 6(x) and cultural with y=7(x) and x was the number of hours in each but was worng